OpenStudy (anonymous):

Choose the correct description of the graph of the compound inequality x - 3 < -9 or x + 5 greater than or equal to 12 A number line with a closed circle on -6, shading to the left, and an open circle on 7, shading to the right. A number line with an open circle on -6, a closed circle on 7, and shading in between. A number line with an open circle on -6, shading to the left, and a closed circle on 7, shading to the right. A number line with a closed circle on -6, an open circle on 7, and shading in between.

OpenStudy (texaschic101):

x - 3 < - 9 x - 3 + 3 < - 9 + 3 (add 3 to each side to get x by itself) x < -9 + 3 x < - 6 x + 5 >= 12 x + 5 - 5 >= 12 - 5 (subtract 5 from each side to get x by itself) x >= 12 - 5 x >= 7 so we have x is less then -6 and greater then 7 ANSWER : a number line with an open circle on -6, shaded to the left, and a closed circle on 7, shaded to the right :)

OpenStudy (anonymous):

A number line with an open circle on -6, shading to the left, and a closed circle on 7, shading to the right.